OpenStudy (anonymous):

If a circle's circumference is 15 inches, what is its radius? Leave the symbol π in your answer.

OpenStudy (johnweldon1993):

Well we know the circumference = 2*pi*r If the circumference = 15..this means 15 = 2*pi*r We want to solve for 'r' so lets divide everything by 2pi...this will isolate 'r' on one side 15 r = ----- 2pi

OpenStudy (anonymous):

Thanks !

OpenStudy (johnweldon1993):

No problem!